### v1

Abstract
Fully homomorphic encryption (FHE) is a technique that enables statistical processing and machine learning while protecting data including sensitive information collected by such single board computers (SBCs) on a cloud server. Among FHE schemes, the TFHE scheme is capable of homomorphic NAND operation, and unlike other FHE schemes, it can perform any operation, such as minimum, maximum, and comparison operations. However, TFHE requires Torus Learning With Error (TLWE) encryption, which encrypts one bit at a time, resulting in less efficient encryption and larger ciphertext size than the other schemes. In addition, SBCs have a limited number of hardware accelerators compared to servers, making it difficult to perform the same optimization as servers. In this study, we propose a novel SBC-specific design TFHE-SBC to accelerate the client-side TFHE operations and achieve communication and energy efficiency. Experimental results show that the TFHE-SBC encryption is up to 2486 times faster, communication efficiency improves 512 times higher, and 12 to 2004 times more energy efficiency than the state-of-the-art.
Keywords:
Homomorphic Encryption Software Optimization IoT
1 Introduction
Single board computers (SBC), such as the Raspberry Pi, can collect various data from temperature, humidity, air pressure, and illumination sensors at low cost and energy efficiency. When sensitive information is included in the data collected by SBC, fully homomorphic encryption (FHE) is useful to securely analyze the SBC data on a cloud server. FHE is a cryptographic technique that enables operations on encrypted data. The main HE schemes are BGV [1], BFV [2], CKKS [3], and TFHE [4], based on the Learning with Errors (LWE) or Ring LWE (RLWE) problem. Each of these schemes supports different types of plaintext (int, float, scalar, vector, etc.) and different homomorphic operations, so it is necessary to select a scheme that is appropriate for the application. Among the main HE schemes, only the TFHE scheme can perform any operation, such as minimum, maximum, and comparison operations. Therefore, the TFHE scheme supports a wide range of applications such as query execution on encrypted databases [5, 6, 7] and encrypted machine learning [8, 9, 10, 11].
However, the computationally heavy nature of homomorphic encryption operations has still largely prevented the FHE from being ported to the IoT domain. For example, a simple TFHE encryption using TFHEpp [12] library requires a minimum of 4.2 MiB in memory allocation when encrypting only 1 KiB with a commonly used parameter set. Acceleration of client-side operations in the TFHE scheme allows data analysts to realize various applications. The goal of our system is to provide a client device (SBC) with the capability of homomorphically encrypting the data collected by the device to facilitate its offload to an untrusted server. A line of recent works provided different designs for the client-side operations of the CKKS scheme [13, 14, 15]. The CKKS scheme supports real number arithmetic, but unlike TFHE, it cannot evaluate arbitrary circuits. What are the challenges of implementing the TFHE scheme at SBC? TFHE applications require TLWE encryption, which encrypts one bit at a time, resulting in inefficient encryption and larger ciphertext size than other schemes. In addition, SBCs have a limited number of hardware accelerators compared to servers, making it difficult to perform the same optimization as servers. To make such a SBC-unfriendly scheme feasible in IoT systems, we list the following three requirements:
  1. R1:
    Computation efficient on SBCs without special accelerators and the server.
  2. R2:
    Communication efficient between SBCs and the server.
  3. R3:
    Energy efficient solution for encryption on SBCs.
Table 1: Benchmark results of client-side operations for the TFHE scheme on Raspberry Pi Zero 2W. TFHE-SBC outperforms other methods in encryption time, communication cost, memory usage, and energy consumption.
Plaintext Methods Encryption Time [ms] Ciphertext [KiB] RAM [KiB] Energy [mJ]
256 bit TLWE by TFHEpp 1922.6 620 $\times$ 1024 128 $\times$ 1217.5 3.6 $\times$ 2499.4 499.9 $\times$
TRLWE by TFHEpp 46.1 14.9 $\times$ 8 1 $\times$ 332.2 0.99 $\times$ 64.5 12.9 $\times$
Ours 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
1024 bit TLWE by TFHEpp 7705.6 2486 $\times$ 4096 512 $\times$ 4297.7 12.8 $\times$ 10017.3 2003.5 $\times$
TRLWE by TFHEpp 46.1 14.9 $\times$ 8 1 $\times$ 332.2 0.99 $\times$ 64.5 12.9 $\times$
Ours 3.1 1 $\times$ 8 1 $\times$ 336.2 1 $\times$ 5.0 1 $\times$
4096 bit TLWE by TFHEpp 30781.2 2386 $\times$ 16384 512 $\times$ 16599.04 46.1 $\times$ 40015.6 1914.6 $\times$
TRLWE by TFHEpp 183.7 14.2 $\times$ 32 1 $\times$ 356.2 0.99 $\times$ 257.1 12.3 $\times$
Ours 12.9 1 $\times$ 32 1 $\times$ 360.2 1 $\times$ 20.9 1 $\times$
8192 bit TLWE by TFHEpp 61587.5 2406 $\times$ 32768 512 $\times$ 33003.5 84.1 $\times$ 80063.8 1929.2 $\times$
TRLWE by TFHEpp 370.8 14.5$\times$ 64 1 $\times$ 388.2 0.99 $\times$ 519.1 12.5 $\times$
Ours 25.6 1 $\times$ 64 1 $\times$ 392.2 1 $\times$ 41.5 1 $\times$
To address the above requirements, we propose TFHE-SBC, the first TFHE framework targeted for SBCs that can simultaneously achieve efficient encryption and communication. Existing implementations of the TFHE scheme use inefficient TLWE encryption, but TFHE-SBC differs from existing implementations in that it uses TRLWE encryption, which is optimized for SBCs. As shown in Figure 1, it consists of a device library that encrypts multiple bits at once using TRLWE encryption on the client side and an adapter library that converts them into the TLWE ciphertext needed for the TFHE application on the server side. We also propose a modified random number generation algorithm and the memory reuse strategy to improve the efficiency of TRLWE encryption on the limited computational resources of the SBC. The main contributions of this work are summarized as follows.
  • To the best of our knowledge, we first employ ciphertext type switching, which converts TLWE to TRLWE, to achieve efficient encryption and communication efficiency on SBCs. This solution also improves energy efficiency.
  • We identify the specific memory and performance challenges related to enabling TFHE on SBCs and discuss several techniques to overcome them. In particular, we show how to design sampling algorithms without environment-dependent optimization. We further investigated the feasibility of accelerating polynomial multiplication with the VideoCore IV (VC4) GPU installed in the Raspberry Pi Zero 2W and experimentally showed that there are accuracy issues.
  • We provide encryption time, communication costs, memory usage, and energy consumption results for a variety of configurations enabled by TFHE-SBC for a parameter set that enables insightful analytics. As sketched in Table 1, we achieved 15 to 2486 times faster encryption, up to 512 times more efficient communication, up to 12.8 times more memory efficiency, and 12 to 2004 times more energy efficiency than existing implementations.
Figure 1: End-to-end TFHE deployment solution by TFHE-SBC.
2 Preliminaries
Notations. A few mathematical objects will be needed to understand this work. $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$ the ring of integer polynomials modulo the cyclotomic polynomial $X^{N}+1$, with power of 2. $\mathcal{R}_{q}=(\mathbb{Z}/q\mathbb{Z})[X]/(X^{N}+1)$, i.e., the same ring of integers $\mathcal{R}$ as above, but this time the coefficients are modulo $q$. Observe that we often note $\mathbb{Z}/q\mathbb{Z}$ as $\mathbb{Z}_{q}$. The letter “T” in TFHE [4] refers to the real torus $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$. Basically, $\mathbb{T}$ is the set $[0,1)$ of real numbers modulo $1$.
2.1 General LWE
The security of the scheme is based on a lattice problem called LWE and its variants, such as RLWE. BGV, BFV, CKKS, and TFHE schemes are based on RLWE or LWE ciphertexts. In this section, we will use a generalization that includes both of them, called General LWE, or GLWE in short.
Let $p$ and $q$ be two positive integers, such that $p\leq q$, and let us define $\Delta=q/p$. In TFHE, $p$ and $q$ are often chosen to be powers of two: if they are not, a rounding at the moment of encoding of messages should be applied. We will call $q$ ciphertext modulus, $p$ plaintext modulus, and $\Delta$ scaling factor.
Key Generation. We sample the secret key $\vec{S}=(S_{0},\ldots,S_{k-1})\in\mathcal{R}^{k}$ with uniform binary distribution as $k$ polynomials of degree smaller than $N$.
Encryption. A GLWE ciphertext encrypting the message $M$ under the secret key $\vec{S}$ is a tuple:
$(A_{0},\ldots,A_{k-1},B)\in GLWE_{\vec{S},\sigma}(\Delta M)\subseteq\mathcal{R}_{q}^{k+1}$
$\text{where}\;B=\sum_{i=0}^{k-1}A_{i}\cdot S_{i}+\Delta M+E\in\mathcal{R}_{q}$
To encrypt the message $M\in\mathcal{R}_{p}$, we need to sample a uniformly random mask $A_{i}\in\mathcal{R}_{q}$ and $E\in\mathcal{R}_{q}$ which has coefficients sampled from a Gaussian distribution $\chi_{\sigma}$.
Since FHE encrypts by adding noise to plaintext, noise accumulates in the ciphertext as homomorphic operations are repeated. The ciphertext cannot be decrypted to the correct value if the noise exceeds a threshold. There is a special type of operation called bootstrapping [16], which reduces the noise of a ciphertext.
Construction of LWE and RLWE. When we instantiate GLWE with $k=n\in\mathbb{Z}$ and $N=1$, we get LWE. When we instantiate GLWE with $k=1$ and $N$ as a power of 2, we get RLWE. LWE differs in that the message is a scalar, while RLWE differs in that the message is a polynomial. Moreover, LWE encryption contains the inner product of the uniformly random mask and the secret key. On the other hand, RLWE encryption contains the polynomial multiplication of the polynomial $A$ and the secret key $S$.
2.2 TFHE Scheme
The TFHE scheme handles three forms of ciphertext: TLWE, TRLWE, and Torus Ring Gentry Sahai Waters (TRGSW). TLWE and TRLWE are a construction of LWE and RLWE over the torus, which is a set of fractional parts of real numbers, respectively. A TLWE ciphertext represents a Boolean value. In contrast, TRLWE represents a vector of Boolean values of length $N$. TRGSW contains multiple TRLWE ciphertexts and is used to represent the bootstrapping key required for the bootstrapping process in the TFHE scheme. The basic ciphertext format is TLWE.
Sample Extraction. Sample Extraction [4] is a component of the bootstrapping process of the TFHE scheme. This operation takes as input a TRLWE ciphertext which encrypts a polynomial message and extracts the encryption of one of the coefficients of the message as TLWE ciphertext. The operation does not increase the noise and simply consists of “copy-pasting” some of the coefficients of the TRLWE ciphertext into the output TLWE ciphertext. The detailed implementation is shown in Algorithm 1.
2.3 Multiplying Polynomials via DWT
Polynomial multiplication is required in (T)RLWE encryption and in the bootstrapping process of the TFHE scheme. In this section, we describe polynomial multiplication by the ordinary discrete Fourier transform (DFT) and polynomial multiplication by the discrete weighted transform (DWT). Let us define polynomials up to degree $N-1$ as $f(x)=f_{0}+f_{1}x\cdots+f_{N-1}x^{N-1}$ and $g(x)=g_{0}+g_{1}x\cdots+g_{N-1}x^{N-1}$. Where $f$ and $g$ are assumed to be periodic such that $f(x+N)=f(x),g(x+N)=g(x)$. The convolution $f*g$ for discrete values is defined as $(f*g)(x)=\sum_{n=0}^{N-1}f(n)g(x-n)$ and is equivalent to polynomial multiplication modulo $x^{N}-1$. The polynomial $F$, which is a DFT of the polynomial $f$, is defined as $F(t)=\sum_{x=0}^{N-1}f(x)e^{-i\frac{2\pi tx}{N}}$. The inverse discrete Fourier transform (IDFT) is defined as $f(x)=\frac{1}{N}\sum_{t=0}^{N-1}F(t)e^{i\frac{2\pi tx}{N}}$. $e^{i\frac{2\pi tx}{N}}$ is called twiddle factors. To obtain $f*g$, we need to compute DFT $\mathcal{F}(f)$ and $\mathcal{F}(g)$, multiply each coefficient of $\mathcal{F}(f)$ and $\mathcal{F}(g)$, and perform IDFT. The computational complexity of the DFT is $\mathcal{O}(N^{2})$, but for $N=2^{k}$, the computational complexity of the polynomial multiplication can be reduced to $\mathcal{O}(N\log N)$ by the fast Fourier transform (FFT) and inverse fast Fourier transform (IFFT). The DWT is needed since TRLWE requires polynomial multiplication modulo $x^{N}+1$, i.e., negacyclic convolution. The polynomial after DWT is $F^{\prime}(t)$, and DFT/IDFT is performed by multiplying it by the weight $w_{x}=e^{i\frac{2\pi tx}{2N}}$ as follows.
$F^{\prime}(t)=\sum_{x=0}^{N-1}w_{x}f(x)e^{-i\frac{2\pi tx}{N}},\quad f(x)=\frac{1}{w_{x}N}\sum_{t=0}^{N-1}F^{\prime}(t)e^{i\frac{2\pi tx}{N}}$
3 Related Work
In this section, we present research on software and hardware implementations of client-side FHE operations (See Table 2 for a summary) and Transciphering, which reduces the amount of communication between client and server.
SEAL-Embedded [15] has accelerated CKKS encoding and encryption processes on embedded devices such as Azure Sphere Cortex-A7, Nordic nRF52840 Cortex-M4. The client-side operations of the CKKS scheme consist of encoding and RLWE encryption processes, requiring the implementation of FFT for encoding and polynomial multiplication and random number generation for RLWE encryption. In SEAL-Embedded, each layer is optimized to implement a different random number generation algorithm, faster polynomial multiplication using Number Theoretic Transform (NTT), and memory reuse. TFHE-SBC is the first TFHE framework that enables a client to execute TFHE encryption efficiently on SBCs.
[13, 17, 18, 14] are FPGA-based accelerators for FHE. RACE [19] and RISE [20] implement encryption of CKKS schemes, including NTT on ASICs, but the encoding process before encryption is implemented in software. Aloha-HE [14] implements a faster encoding and encryption process for the CKKS scheme on FPGA. The implementation includes an FFT unit for encoding, an NTT unit for encryption, and a random number generator unit. The encoding and decoding speeds of CAEA [13] are both 2.01 times faster, and the encryption and decryption speeds of CAEA are 1.13 times and 3.04 times faster than Aloha-HE, respectively. However, ASIC implementation is less flexible than software implementation in responding to changing requirements. Furthermore, the previous work is limited to the CKKS scheme, and client-side operations acceleration for the TFHE scheme, which can support flexible applications, has not yet been implemented.
Transciphering [21, 22, 23, 24, 25] allows for lower communication overload between the client and the server compared to using any homomorphic encryption scheme alone. A client does not need to encrypt all its messages using an HE algorithm (except the symmetric key, such as the AES key); all the messages can be encrypted using only a symmetric cipher. The server obtains the message encrypted with HE by homomorphically evaluating the AES decryption process. On the other hand, this conversion on the server requires at least about 6 seconds [21] under the BGV scheme and about 9 seconds [23] under the TFHE scheme. Therefore, although Transciphering reduces the amount of client-server communication, it tends to increase the overall latency of the application. In our framework, the load on the server caused by reduced communication costs is much less than that of Transciphering.
Table 2: Related work on homomorphic encryption on the client.
  [13] [17] [18] [19] [20]   [14] [15]  Ours
Support TFHE scheme
Software acceleration 1,2 1 1 2 2 1
Consideration of communication costs
Optimization of random number generation
Reporting power consumption
  • 1
    Accelerating on FPGA platforms.
  • 2
    Accelerating on ASIC platforms.
4 Speeding up TLWE Encryption via Sample Extraction
In the TFHE scheme, the application is executed by performing logical operations on TLWE-encrypted data. Therefore, the client device must perform TLWE encryption on a Boolean value. In this section, we present the challenges of TLWE encryption in terms of computational complexity and ciphertext size and describe the proposed protocol. The parameters to be set for the TFHE scheme are the values of Table 3 that satisfy 128-bit security.
Table 3: Parameters achieve 128-bit security in the TFHE scheme.
$q$ $2^{32}$ The modulus for discretizing Torus.
$N$ $2^{10}$ The length of the TLWE ciphertext. The dimension of TRLWE ciphertext.
$\sigma$ $2^{-25}$ The standard deviation of the noise for the fresh TLWE and TRLWE ciphertext.
4.1 Threat Model and Assumptions
For our threat model, we consider a client or user to be a trusted party capable of performing key generation, encryption, and decryption in trusted environment, and a server to be an untrusted party to which the client would like to offload computation on sensitive data. We discuss the threat protections that our system is able to provide at a high level and our deployment assumptions.
First, we assume that any transfer of secret keys for TFHE can be transferred to the device securely, either before device deployment or over a secure channel. We also assume that the deployed IoT device has access to standard network protection mechanisms for confidentiality, integrity, and freshness of transmitted data. These network security mechanisms are still required on top of homomorphically encrypted ciphertexts since they are necessary for protection against active attacks and secure entire network messages (including any network packet headers) rather than just the ciphertext data itself. In our implementation of TFHE-SBC on Raspberry Pi Zero 2W, for example, communication channels are protected by the TLS protocol.
Furthermore, since no party can learn anything about the underlying data through the encrypted ciphertexts alone, HE is able to provide protection against all aforementioned cloud attackers. HE itself is not robust against malicious cloud attackers, who can manipulate encrypted ciphertexts in the cloud in addition to viewing server state. However, HE can still offer data privacy against these adversaries as well. A remaining threat that we consider is the threat of direct attacks on the device. We consider these threats largely out of scope, since we assume that an attack on the device can easily access the underlying private user data (e.g., sensor samples) or the source of data itself. Since an attacker would already have access to the raw data in this scenario, any additional protections on ciphertexts, keys, or error polynomials would be extraneous. However, if a secret key is stored on the device, a device attacker could theoretically also obtain and use this key to decrypt all previously encrypted information stored in the cloud, thereby expanding the scope of a device attack to a potentially large amount of previously offloaded user data. To mitigate this situation, users can choose to update their device secret keys frequently, thus limiting the interval of data that a device attacker could decrypt using the user’s secret key. This would mean that any data encrypted under the old secret key would no longer be able to be analyzed alongside any data encrypted with the new secret key. However, this is not a problem for private inference applications [8] because they do not store ciphertext in the cloud.
Finally, we note that the post-quantum nature of HE additionally provides some robustness against quantum attacks. In particular, attacks mounted by a quantum-time adversary on HE ciphertexts would not be able to access the underlying private data. For our implementation, we leave quantum attacks out of scope. However, if quantum protection is desired, several existing libraries [26, 27] offer solutions for quantum-secure encryption of messages across an untrusted channel (such as a public network).
4.2 Costs Analysis and Our Proposal
Considering the difference between TLWE and TRLWE configurations, we analyze the computational cost of encryption and the communication cost, i.e., ciphertext size.
Computation costs. TLWE encrypts plaintext per 1-bit and includes $N$ times uniformly distributed random number generation, the inner product of a vector of length $N$, and one time normally distributed random number generation.
TRLWE encryption involves $N$ bit encryption, $N$ times uniformly distributed random number generation, $N$ length polynomial multiplication, and $N$ times normally distributed random number generation.
Therefore, the time complexity of TLWE encryption is $\mathcal{O}(N)$ per bit, and that of TRLWE encryption is $\mathcal{O}(\log N)$ per bit since the time complexity of polynomial multiplication using FFT is $\mathcal{O}(N\log N)$.
Communication costs. From $(a_{0},\ldots,a_{N-1},b)\in LWE_{\vec{s},\sigma}(\Delta m)\subseteq\mathbb{Z}_{q}^{N+1}$, the TLWE ciphertext requires an array with length $N+1$ per bit. From $(A,B)\in RLWE_{{S},\sigma}(\Delta M)\subseteq\mathcal{R}_{q}^{2}$, the TRLWE ciphertext requires an array of length $2N$ per $N$ bit, so the size of the TLWE ciphertext is $\frac{N+1}{2}$ times larger than that of TRLWE ciphertext.
Replace TLWE with TRLWE. It is clear that TLWE encryption is computationally more expensive and has a larger ciphertext size than TRLWE encryption. In order to avoid TLWE encryption at the client, we propose to replace TLWE encryption with TRLWE encryption and Sample Extraction [4]. As described in Section 2.2, Sample Extraction converts a single TRLWE ciphertext into $N$ TLWE ciphertexts. In general, Sample Extraction is a component used in the bootstrapping process and is not intended to reduce client-side computational and communication costs. The client sends TRLWE ciphertext to the server, and the server converts it into TLWE ciphertext using Sample Extraction, thereby reducing the client’s execution time and ciphertext size. Since Sample Extraction does not increase the noise, this protocol is also superior to Transciphering [23, 24, 25] that involves the bootstrapping. The protocol of the proposed method is shown below.
  1. 1.
    The trusted party generates the private key, Bootstrapping key, and key switching key, and shares the private key with the client device.
  2. 2.
    The client device encrypts the message with TRLWE and shares the ciphertext with the cloud server. TRLWE encrypts by packing multiple bits together.
  3. 3.
    The cloud server converts TRLWE ciphertext to TLWE ciphertext using Sample Extraction.
  4. 4.
    The cloud server executes the application by homomorphic operations and shares the computation results with the trusted party.
  5. 5.
    The trusted party decrypts the result.
5 Bottleneck Analysis in TRLWE Encryption
In this section, we analyze the bottleneck of TRLWE encryption in TFHEpp, to identify areas for optimization of TRLWE encryption implemented in TFHE-SBC. Among the OpenFHE [28] and TFHEpp [12] libraries that handle TFHE schemes, the TFHEpp library has more functions than the others [29]. Moreover, TFHEpp is considered faster than implementation [30] by the authors of the original TFHE, and our work is the first TFHE framework that enables a client to execute TFHE encryption efficiently on single-board computers. Therefore, we employ the TFHEpp implementation as our baseline.
As described in Section 2.1, TRLWE encryption consists of pseudorandom number generation sampled from the uniform and Gaussian distribution, and polynomial multiplication. Table 4 shows the breakdown of TRLWE encryption on the Raspberry Pi Zero 2W using trlweSymEncrypt function in TFHEpp. “Uniform Sampling” and “Gaussian Sampling” refer to random number generation from their respective distributions, and “PolyMul” refers to polynomial multiplication. The most time-consuming operation is the generation of Gaussian random numbers, which account for 83.1% of the total execution time. Since random numbers sampled from the uniform and Gaussian distribution are generated the same number of times, random numbers sampled from the Gaussian distribution are generated slower than those sampled from the uniform distribution per random number generation. PolyMul has less impact on encryption time than random number generation. We give details about the bottleneck of each component and further discuss their optimization.
5.1 Hardware Dependent Pseudorandom Number Generator
TFHEpp can use /dev/urandom provided by the Linux OS or Randen [^1] https://github.com/google/randen as a cryptographically secure pseudorandom number generator (CSPRNG). Randen is an implementation that requires an AES accelerator, but at this time, AES accelerators are not supported except for Raspberry Pi 5, so Table 4 is the execution time when using /dev/urandom. Random numbers obtained from CSPRNG are used to generate uniform and Gaussian random numbers, so the speed of pseudo-random number generation affects the speed of sampling from uniform and Gaussian distribution. Moreover, it has been reported that random number generation using /dev/urandom slows down the encryption speed [31]. [31] pointed out that TRGSW encryption is 35 times faster on the SBC, which supports AES accelerators, than on the Raspberry Pi 4. It is necessary to implement a CSPRNG that is fast even in environments where AES accelerators are not supported like the Raspberry Pi Zero series.
5.2 Gaussian Random Number Generator
The Gaussian random number generation algorithms include Box-Muller method [32], polar method [33], Kinderman method [34], Monty-Python method [35], and Ziggurat method [36]. In TFHEpp, Gaussian random numbers are generated by the std::normal_distribution function included in the C++ standard library, and the random number generation algorithm is the polar method. The polar method eliminates the trigonometric calculations required by the Box-Muller method but requires resampling with a probability of about 21.5%. This means that the expected number of uniformly distributed random numbers generated for generating a single Gaussian random number is 1.27. The ziggurat method, which has a lower rejection probability than the polar method, can speed up the process.
Table 4: Breakdown of trlweSymEncrypt function in TFHEpp. Most of the non-optimized TRLWE encryption time consists of random number generation.
 Uniform Sampling  Gaussian Sampling  PolyMul  Total
7.7 [ms] 38.3 [ms] 0.16 [ms] 46.1 [ms]
16.7 % 83.1 % 0.35 % 100 %
5.3 Multiplying Polynomials and Applicability of VC4 GPU
For the polynomial multiplication, we enabled the option in TFHEpp to use the implementation for AArch64 of the FFT library SPQLIOS [^2] https://github.com/tfhe/experimental-tfhe/tree/master/circuit-Bootstrapping/src/spqlios. SPQLIOS does not use a real FFT, but rather packs the imaginary part of the input as a complex number, so that a polynomial of length $N$ can be transformed using an FFT with a polynomial of length $\frac{N}{2}$ as input. We observed that this SPQLIOUS optimizes well for CPU-only implementations. Does GPU on Raspberry Pi enable faster polynomial multiplication?
Broadcom VideoCore IV [^3] https://docs.broadcom.com/doc/12358545 is the GPU on the Raspberry Pi. Quad Processing Unit (QPU), the main arithmetic unit of the VC4, has four floating-point arithmetic units and operates as a 16-parallel SIMD over four cycles. The floating-point arithmetic units can simultaneously perform addition and multiplication in a single cycle. Four QPUs make up a slice, and there are three slices in VC4. Each of them operates at 0.25GHz, giving a peak performance of 24GFlops. VC4 uses the host memory and reads and writes by DMA. Of the 512 MB of host memory, VC4 is allocated 128 MB by default. We explore whether the FFT, a component of polynomial multiplication, can be accelerated by VC4.
Table 5: FFT/IFFT Execution and PolyMul Accuracy on VC4
 FFT [ms]  IFFT [ms]  RRMSE
CPU 0.054 0.061 $4.28\times 10^{-7}$
VC4 (GPU) 0.034 0.036 0.015
The FFT implementation GPU_FFT [37] provided by Raspberry Pi. This provides all power-of-2 FFT lengths between $2^{8}$ and $2^{22}$ points inclusive. GPU_FFT uses a kernel character device known as the “mailbox” for communication between the ARM and the Videocore. The channels include memory allocation, code execution, the getting temperature, firmware revision, etc. For parallelism, VC4 has 12 QPUs; but GPU_FFT only uses 8. It happens that bit-rotated writes can be done very neatly through the Vertex Pipe Memory and DMA controller VDW by 8 QPUs working in parallel.
Execution time. Ignoring the memory copy between the CPU and VC4, Table 5 shows that the FFT and IFFT are executed 1.59 times to 1.69 times faster on the VC4 than on the CPU by SPQLIOS with $N=2^{10}$. Copying buffer from CPU to GPU requires 0.0045 ms, and from GPU to CPU requires 0.00096 ms, with the same $N$. The memory transport time between the CPU and GPU does not exceed the main process, and the GPU provides a speed-up effect.
Accuracy. PolyMul by FFT is known to have a small error due to floating-point arithmetic, but FFT on VC4, which supports only single precision, has additional errors. In the TFHE scheme, since the maximum number of coefficients for PolyMul before rounding to the modulus $q$ is $N*(2^{31}-1)$, the number of bits required for exact PolyMul is $4231+\log_{2}N=$ 42 bits, where $N=2^{10}$. The CPU implementation can retain this maximum value because it can operate in double precision. However, GPU_FFT uses single-precision floats for data and twiddle factors, which does not achieve sufficient PolyMul accuracy. This is because the QPU in VC4 contains two single-precision floating-point units, one for addition and one for multiplication, each operating on a vector of four single-precision numbers. We compare the relative root mean square error (RRMSE) between native PolyMul implementation and PolyMul with FFT / IFFT, defined as $\text{RRMSE}(X,Y)={\sqrt{\frac{1}{n}\sum_{i=1}^{n}|X_{i}-Y_{i}|^{2}}}/{\sqrt{\sum_{i=1}^{n}|X_{i}|^{2}}}$. The native implementation provides exact polynomial multiplication, but its computational cost is $\mathcal{O}(N^{2})$. From Table 5, PolyMul on the single-precision VC4, the RRMSE is closer to 1, which is more inaccurate than PolyMul on CPU. Therefore, we execute PolyMul entirely on the CPU to preserve accuracy, although FFT/IFFT on VC4 is faster.
6 TFHE-SBC
This section describes TFHE-SBC, which streamlines the client-side operations of the TFHE scheme in SBCs based on the bottleneck analysis in Section 5. TFHE-SBC consists of the device library and adapter library, using an implementation derived in part from the TFHEpp library. The device library contains optimized functions for TRLWE encryption and is installed on the client side. The adapter library converts the data encrypted in the device library into a format that can be handled by the TFHEpp library and is installed on the server side. An overview of the implementation is shown in Figure 2.
Figure 2: Composition of TFHE-SBC: The SBC performs TRLWE encryption, and the server converts the TRLWE ciphertext into TLWE ciphertext. FFT for the secret key is performed only once and reused for subsequent encryption.
6.1 Device Library
This section presents the structure of the device library, including optimized random number generation and PolyMul.
Pseudo-random number generation with BLAKE2. In Section 5.1, we showed that pseudo-random number generation by /dev/urandom is the bottleneck. In TFHE-SBC, only the random seed is taken from /dev/urandom, and the random seed is expanded by BLAKE2444https://www.blake2.net/. BLAKE2 is a cryptographic hash function faster than MD5, SHA-1, SHA-2, and SHA-3, yet is at least as secure as the latest standard SHA-3. This is also employed in the OpenFHE library.
Gaussian random number with Ziggurat method. In Section 5.2, we showed that the generation of Gaussian random numbers by the polar method is the bottleneck of TRLWE encryption in TFHEpp. In TFHE-SBC, we implement the ziggurat method for Gaussian random number generation, which is faster than the polar method. The ziggurat method [36] is considered the fastest of the five algorithms listed in Section 5.2. Belonging to the class of rejection sampling algorithms, it relies on an underlying source of uniformly distributed random numbers, typically from a pseudo-random number generator, as well as precomputed tables. With a probability of 98.8%, random numbers can be generated with two floating-point multiplications, and the expected number of uniformly distributed random numbers when generating one Gaussian random number is 1.01, which is less than that of the polar. In TFHE-SBC, Gaussian random number generation is based on the boost::random::normal_distribution function in the boost library [^5] https://www.boost.org/ to implement the ziggurat method.
Memory reuse in PolyMul. In TFHEpp implementation of TRLWE encryption shown in Section 5, the execution time for polynomial multiplication was 0.35%, but when the random number generation acceleration described above is applied, the polynomial multiplication accounts for a larger percentage of the execution time. As long as GPU-enabled FFT is not effective, we consider cost reduction of PolyMul on CPU. In TFHEpp, the FFT implementation has already been refined by SPQLIOS, TFHE-SBC further reduces computation costs by reusing the secret key after the Fourier transform. In SBC, when the encryption function is called multiple times, the uniform noise must be resampled and Fourier transformed each time, but the secret key is Fourier transformed only once and then reused. This strategy is also illustrated in Figure 2.
Algorithm 1 TRLWEtoTLWEs a.k.a, Sample Extraction [4]
0:  $(A,B)\in\text{TRLWE}(\Delta M)$
0:  $(a_{0,0},\ldots,a_{0,N-1},b_{0})\in\text{TLWE}(\Delta m_{0})$, $\ldots$, $(a_{N-1,0},\ldots,a_{N-1,N-1},b_{N-1})\in\text{TLWE}_{(}\Delta m_{N-1})$
  for $0\leq h<N$ do
     for $0\leq i\leq h$ do
        $a_{h,i}=A[h-i];$
     end for
     for $h+1\leq i<N$ do
        $a_{h,i}=-A[N+h-i];$
     end for
     $b_{h}=B[h];$
  end for
  return  $(a_{0,0},\ldots,a_{0,N-1},b_{0}),$ $\ldots$,… , $(a_{N-1,0},\ldots,a_{N-1,N-1},b_{N-1})$
6.2 Adapter Library
We additionally provide an “adapter” server module to convert data encrypted by TFHE-SBC to be compatible with the TFHEpp library. The adapter library converts TRLWE ciphertexts received from the client into TLWE ciphertexts that can be handled by TFHEpp as shown in Algorithm 1. The conversion process is based on SampleExtractIndex by TFHEpp. This function copies a part of the TRLWE ciphertext array to the TLWE ciphertext. SBC inputs the TRLWE ciphertext $\text{TRLWE}(\Delta M)$ where $M=\sum_{j=0}^{N-1}m_{j}X^{j}\in\mathcal{R}_{p}$, then the adapter library outputs TLWE ciphertexts $\text{TLWE}(\Delta m_{0}),\text{TLWE}(\Delta m_{1}),\ldots,\text{TLWE}(\Delta m_{N-1})$. Ciphertext processed by the adapter library can be passed to any TFHEpp-implemented application.
7 Experiments
As discussed in Section 1, we attempt to optimize client-side TFHE operations and communication costs. This section answers the following questions via extensive empirical evaluations to confirm how much these requirements are satisfied.
  1. Q1:
    How fast is the running time on the SBC and the server?
  2. Q2:
    How large is the ciphertext size exchanged between the SBC and the server?
  3. Q3:
    How does the power consumption of TFHE-SBC in client-side operation compare to other methods?
The parameters for TLWE and TRLWE are the values in Table 3, which are the same as lvl1param in the TFHEpp library. This value achieves 128-bit security. All of our experiments are conducted on Raspberry Pi Zero 2W with an ARM 1GHz 4-core CPU and 512MB RAM as SBC and a machine with an Intel 1.9GHz 56-core CPU and 512GB RAM as the cloud server. We implemented TFHE-SBC using C++17 and complied with GCC-10.2.1. Our implementation is available on https://anonymous.4open.science/r/TFHE-SBC-A178.
For Q1 and Q2, we compare the encryption time on the SBC and ciphertext size of TLWE and TRLWE encryption and show the conversion time from TRLWE ciphertext to TLWE ciphertext on the server. We also compare the running time of TRLWE encryption for each condition of random number generation and PolyMul methods and discuss the optimal combination to be employed in TFHE-SBC. We also report memory usage for each configuration.
For Q3, we connect the UM24 USB meter between the power supply and the device to measure the power consumption.
Table 6: Runtime and data memory requirements for each noise sampling.
Sampling Configuration Runtime [ms] RAM [KiB]
Uniform /dev/urandom 7.71 192.1
BLAKE2 0.92 196.2
Gaussian /dev/urandom 38.26 192.1
Polar
/dev/urandom 15.36 192.2
Ziggurat
BLAKE2 4.98 196.2
Polar
BLAKE2 1.98 196.2
Ziggurat
Table 7: Breakdown of PolyMul
Configuration Runtime [ms] RAM [KiB]
FFT Coeff-Wise Mult IFFT
on-the-fly 0.103 0.0086 0.0611 314.6
reuse 0.065 0.0086 0.0611 314.6
7.1 Experimental Results
To answer Q1, we first evaluate the computational cost on the SBC side and then on the server side. We also show the optimization effects in TRLWE encryption in detail for the uniform and Gaussian noise generation and PolyMul, respectively.
End-to-end computation costs on SBCs. Table 1 compares the execution time, ciphertext size, and energy consumption of TLWE encryption (tlweSymEncrypt function) and TRLWE encryption (tlweSymEncrypt function) by the TFHEpp library as the baseline with TFHE-SBC. The plaintext size is set from 256 bits to 8192 bits. TFHE-SBC library runtime is the value when implemented with the fastest combination of random number generation and polynomial multiplication as described in Section 6.1. TFHEpp in environments that do not support AES accelerators is implemented using /dev/urandom as CSPRNG and the polar method for Gaussian noise. In terms of execution time, TRLWE encryption encrypts $N$ bits at a time, so the encryption times for 256-bit and 1024-bit encryption are the same. The optimized TRLWE with TFHE-SBC is approximately 15 times faster than TRLWE encryption with TFHEpp library and approximately 2490 times faster than TLWE encryption. For RAM consumption, TFHE-SBC is 12.8 times more efficient for 1024-bit encryption than baseline TLWE encryption. For TRLWE, memory usage was 336.2 KiB for our method, 1.01 times the baseline. The increase of 4 KiB memory usage in TFHE-SBC is due to the random number generation being replaced by BLAKE2, as can be observed from Table 6. This means that TFHE-SBC achieved 15 times faster encryption without much increase in memory usage from non-optimized TRLWE.
(a) Execution time
(b) Memory usage
Figure 3: Performance of TRLWEtoTLWEs. It is clear that end-to-end encryption of TFHE-SBC is faster than TLWE encryption on the client, even when conversion time is taken into account.
Switching from TRLWE to TLWE. We evaluate the server-side computational cost of the TFHE-SBC adapter library. Note that if the client is TLWE encrypted, such costs are ignored. Figure 3(a) shows the execution time of TRLWEtoTLWEs on the server side with TFHE-SBC’s adapter library. The execution is single-threaded. For a 1024-bit plaintext, TRLWE encryption with the TFHE-SBC’s device library requires 3.1 ms on the client device and 1.0 ms for conversion on the server, for a total of 4.1 ms. In this case, TLWE encryption requires 7705.6 ms from Table 1. Therefore, end-to-end encryption of TFHE-SBC is faster than TLWE encryption on the client, even when the total execution time of TRLWEtoTLWEs is taken into account. Figure 3(b) shows that the peak memory usage of TRLWEtoTLWE increases linearly. TRLWEtoTLWE requires 8KiB TRLWE ciphertext as input and outputs 4MiB TLWE ciphertexts, so at least 4plaintext $4*\frac{\text{plaintext size}}{N}$ MiB is reserved. Additional memory is required to concatenate the output TLWE by std::copy.
Comparison of the noise sampling. From Table 6, changing from /dev/urandom to BLAKE2 as the CSPRNG had the greatest speedup effect. Since CSPRNG is involved in generating random numbers sampled from uniform and Gaussian distributions, it indicates that an improvement in the speed of pseudorandom number generation affects the speed of sampling from each distribution. BLAKE2 generates random numbers 8 times faster than /dev/urandom for both uniform and Gaussian distributions with an additional 4 KiB increase in memory usage. As shown in Section 6.1 and Table 6, the execution time of Gaussian number generation improved by employing the Ziggurat method rather than the polar method employed in TFHEpp. When implemented in the same CSPRNG, Gaussian sampling using the Ziggurat method is more than twice as fast as Gaussian sampling using the polar method. These results are derived from the difference in rejection probabilities of each method.
Comparison of the polynomial multiplication. We test the performance of PolyMul with two configurations, as shown in Table 7. In the “on-the-fly” version, we execute FFTs for the secret key and the uniform noise, so the FFT is performed twice for each encryption. In the “reuse” flavor, the FFT value of the secret key is precomputed. It requires less computation than the “on-the-fly” case since the FFT for the secret key only needs to be calculated once per encryption. Peak memory usage is the same for both cases. PolyMul “re-use” case, which eliminates one FFT execution, improves the computational cost of the PolyMul “on-the-fly” case by 28%.
To answer Q2, Table 1 shows the communication costs between SBCs and the server. The communication cost here refers to the ciphertext size. The ciphertext size of TLWE, which encrypts one bit at a time, is approximately 512 times larger than that of TRLWE, which encrypts $N$ bits at once. This result is consistent with the theoretical result in Section 4.2. That is, generating TRLWE ciphertext leads to a reduction in the amount of client-server communication.
To answer Q3, we conduct measurements of the SBC’s power consumption. The rightmost column of Table 1 shows the energy consumed in the end-to-end encryption for each plaintext size on SBC. The idle power consumption in Raspberry Pi Zero 2W is 600 mW. TFHE-SBC consumes 5 mJ per encryption, nearly 2000 times less energy than TLWE encryption. In addition, our approach achieves 12 times more energy efficiency than baseline TRLWE encryption. This is because TFHE-SBC has a better encryption runtime.
8 Limitations and Conclusions
We proposed TFHE-SBC, the first TFHE scheme implementation of the client-side operations for SBCs. TFHE-SBC archive the efficient TLWE encryption and reduce the ciphertext size through the TRLWE encryption on SBCs and its conversion to TLWE ciphertext by the server. Furthermore, we built efficient TRLWE encryption with the limited computational resources of SBC and investigated optimal random number generation and polynomial multiplication, including the memory reuse strategy. We also discussed the applicability of GPUs on Raspberry Pi in the TFHE scheme and showed the lack of accuracy of polynomial multiplication. Experimental results show that TFHE-SBC achieved 15 to 2486 times faster encryption, 512 times more efficient communication, up to 12.8 times more memory efficiency, and 12 to 2004 times more energy efficiency than the state-of-the-art on Raspberry Pi Zero 2W. Note that our framework only supports private key encryption, and public key encryption is not implemented yet. It is an interesting challenge to efficiently implement encryption on SBCs that requires a larger number of noise samplings and a much larger public key than the secret key.